A rectangular window that is 1.2 m wide and 1.8 m high has a thickness of 6.2 mm, a thermal conductivity value of 0.27 W/m/°C. The temperature inside the home is 21°C and the temperature outside the home is -4°C. Calculate the rate of heat transfer. 


1
Expert's answer
2021-01-19T07:08:38-0500

The Fourier's law says

"P=\kappa \frac{\Delta T}{\Delta x}A"

"P=0.27\times\frac{21-(-4)}{0.0062}\times 1.2\times 1.8=2352\;\rm W"